OpenStudy (anonymous):

Could someone please tell me how to solve this equation? The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is

OpenStudy (anonymous):

c = 9/5 (f - 32) 5/9 c = (f - 32) 5/9 c +32 =F now put c = 2 in abv relation you get F = 33.33

OpenStudy (anonymous):

you have typed a wrong equation